OpenStudy (anonymous):

Express 5√3x^3 - 2√27x^3 in simplest radical form

OpenStudy (watchmath):

\(5\sqrt{x^2\cdot 3x}-2\sqrt{9x^2\cdot 3x}\) \(5x\sqrt{3x}-6x\sqrt{3x}\) \(-x\sqrt{3x}\)

OpenStudy (anonymous):

64\[\sqrt{5+25}\]